Calculate the variance of an investment with five year returns of 10%, (15%), 12%, 8% and (5%).
Answer #1

Average return=Total return/Total time period

=(10-15+12+8-5)/5=2%

Return (Return-Average return)^2
10 (10-2)^2=64
-15 (-15-2)^2=289
12 (12-2)^2=100
8 (8-2)^2=36
-5 (-5-2)^2=49
Total=538%

Standard deviation=[Total (Return-Average return)^2/(Time period-1)]^(1/2)

=[538/(5-1)]^(1/2)

=[538/4]^(1/2)

=11.60%(Approx).

Variance=Standard deviation^2

=134.5%